OpenStudy (anonymous):

A circle is centered at the point (5, -4) and passes through the point (-3, 2).

OpenStudy (anonymous):

The equation of this circle in standard form is (x +____)2 + (y +____)2 =_____

OpenStudy (anonymous):

@iGreen @welshfella @Girlamerican16

OpenStudy (anonymous):

@shifuyanli

OpenStudy (anonymous):

sorry I haven't really got to this yet, I hope you understand :)

OpenStudy (anonymous):

its fine

OpenStudy (anonymous):

The equation of a circle centered at \((x_0, y_0)\) with radius \(r\): \[(x - x_0)^2 + (y- y_0)^2 = r^2\] Use this equation and use the distance formula to solve for the radius: \[d = \sqrt{(x_1 - x_0)^2 + (y_1 - y_0)^2}\]

OpenStudy (anonymous):

oh ok thanks

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

anybody?

OpenStudy (shamim):

Formule for a circle is (x-h)^2+(y-k)^2=r^2

OpenStudy (shamim):

Given h=5 k=-4

OpenStudy (shamim):

r=radius of ur circle

OpenStudy (shamim):

r=distance frm centre of a circle to any point on the circle

OpenStudy (shamim):

So r=distance frm (5,-4) to the point (-3,2)=?

OpenStudy (anonymous):

i still dont get it

OpenStudy (shamim):

Ok. Can u tell me r=?

OpenStudy (anonymous):

what do you mean?

OpenStudy (shamim):

r=radius of ur circle

OpenStudy (shamim):

Now u hv to know the formula to calculate the distance

OpenStudy (anonymous):

\[r=C/2\]

OpenStudy (anonymous):

with pi

OpenStudy (shamim):

Ok. r=root((5--3)^2+(-4-2)^2)=?